# Concentration of Measure for Distributions Generated via Diffusion Models

### v2

@@ Abstract
We show via a combination of mathematical arguments and empirical evidence that data distributions sampled from diffusion models satisfy a Concentration of Measure Property saying that any Lipschitz $1$-dimensional projection of a random vector is not too far from its mean with high probability. This implies that such models are quite restrictive and gives an explanation for a fact previously observed in the literature that conventional diffusion models cannot capture "heavy-tailed" data (i.e. data $\mathbf{x}$ for which the norm $\|\mathbf{x}\|_{2}$ does not possess a sub-Gaussian tail) well. We then proceed to train a generalized linear model using stochastic gradient descent (SGD) on the diffusion-generated data for a multiclass classification task and observe empirically that a Gaussian universality result holds for the test error. In other words, the test error depends only on the first and second order statistics of the diffusion-generated data in the linear setting. Results of such forms are desirable because they allow one to assume the data itself is Gaussian for analyzing performance of the trained classifier. Finally, we note that current approaches to proving universality do not apply to this case as the covariance matrices of the data tend to have vanishing minimum singular values for the diffusion-generated data, while the current proofs assume that this is not the case (see Subsection 3.4 for more details). This leaves extending previous mathematical universality results as an intriguing open question.
@@ 1 Introduction
One of the most astonishing contributions of deep learning is the advent of generative models for image and video generation. Diffusion-based generative models Sohl-Dickstein et al. [2015], Song and Ermon [2019],Ho et al. [2020],Song et al. [2020],Dhariwal and Nichol [2021],Song et al. [2021],Kingma et al. [2021], Karras et al. [2022], in particular, have enjoyed tremendous success in vision [LDM Rombach et al. [2022], audio [Diffwave Kong et al. [2020]] and text generation [D3PM Austin et al. [2021]]. For an overview of diffusion models and their applications, we refer to the surveys Croitoru et al. [2023] and Yang et al. [2023].
Despite significant progress in training methods, network architecture design, and hyperparameter tuning, there has been relatively little work done on understanding rigorous mathematical properties of the data generated by diffusion models. Through theory and experiments, we argue that images generated by conventional diffusion models are mathematically tractable. In fact, we argue that when the reverse process is a contraction, one can establish a concentration of measure phenomenon for the distribution of the output. Since the latter suggests a form of Gaussian Universality may hold, we compare the generalization error of linear models trained on diffusion data to the generalization error of linear models trained on Gaussian Mixtures with matching means and covariances and observe a close match.
@@
  • We show that distributions that can be generated via diffusion models are far from arbitrary and share many properties with Gaussian distributions in a precise mathematical sense. In particular, this implies the empirical observation made in Pandey et al. [2024] that traditional diffusion models are ineffective for generating distributions with heavy tails, but this could be remedied by training the denoiser on heavy-tailed data and initiating the sampling process with heavy-tailed noise.
  • As discussed in Goldblum et al. [2023] and Nakkiran [2021], one of the major factors that hinders us from having a solid theory of deep learning is the lack of practical assumptions amenable to clean mathematical formulations and analyses regarding the true distributions of data. We believe that our work sheds light on this question in the context of image classification tasks. To elaborate further on this point, note that most image data sets encountered in practice can be approximated well using data sampled from diffusion models. This suggests that in a sense it is sufficient to explain generalization and performance of models for data coming from GMMs. The latter has been a topic of active research recently; see, e.g. Thrampoulidis et al. [2020], Loureiro et al. [2021b].
@@ 2 Related Works
  • In an emerging line of work, many papers have analyzed the output distributions and convergence of diffusion models through the lens of Langevin dynamics. Chen et al. [2022] show that with denoising diffusion probabilistic models (DDPM) and critically damped Langevin Dynamics (CLD), one can efficiently sample from any arbitrary distribution, assuming accurate score estimates - an assumption central to many works in this area. In fact, this work was among the first to provide quantitative polynomial bounds on convergence. Unfortunately, given the high-dimensional nature of the problem, estimating the score function may be practically impossible. Furthermore, it is infeasible to verify the validity of this assumption, as we do not have access to the true score function. And, as evident from the bounds of Mousavi-Hosseini et al. [2023], generating heavy-tailed distributions using Langevin dynamics initialized from the Gaussian distribution is intractable in practice as one needs to run the Langevin dynamics for an exponential number of steps. We refer to Li et al. [2024a] for a brief overview of the existing works on the convergence theory of diffusion models.
@@
  • Seddik et al. [2020] show a form of equivalence between representations generated from Generative Adversarial Networks (GANs) and from GMMs. They considered the Gram matrix of pre-trained classifier representations of the GAN-generated images and show that asymptotically, it possesses the same distribution of eigenvalues as the Gram matrix of Gaussian samples with matching first and second moments.
@@
  • Loureiro et al. [2021b] investigated the generalization error of linear models for binary classification with logistic loss and $\ell_{2}$ regularization. On MNIST and Fashion MNIST, they observed a close match between the real images and the corresponding GMM for the linear models and in the feature map of a two-layer network. Loureiro et al. [2021a] considered a student-teacher model and verified universality for the aforementioned datasets via kernel ridge regression. They also explored the output of a deep convolutional GAN (dcGAN), labeling it using a three-layer teacher network. Using logistic regression for classification illustrated a close match with GMMs on GAN-generated data, but also a deviation from real CIFAR10 images. Goldt et al. [2022] conducted a similar set of experiments and analyzed the generalization error of the Random Features logistic regression using the Gaussian Equivalence property. Furthermore, they considered a dcGAN trained on CIFAR100 and corroborated their theoretical findings on the dataset of images generated by the dcGAN. Then Pesce et al. [2023] studied the student-teacher model for classification and empirically demonstrated the universality of the double descent phenomenon for MNIST and Fashion MNIST. They preprocessed these datasets using a random feature map, with labels generated by a random teacher, for the ridge regression and logistic classification tasks. However, they also observed that the universality of the test error fails to hold while using CIFAR10 without preprocessing with either random feature maps, wavelet scattering, or Hadamard orthogonal projection. Moreover, Dandi et al. [2024] observed that the data distributions generated by conditional GANs trained on Fashion MNIST exhibit Gaussian universality of the test error for generalized linear models. And Gerace et al. [2024] considered mixture distributions with random labels and demonstrated universality of test error of the generalized linear models. The universality part of our work can be considered as an exploration of the same phenomenon for conditional diffusion models trained on significantly larger image datasets.
@@
  • Refinetti et al. [2023] show that SGD learns higher moments of the data as the training continues which exhibited nonuniversality of the test error with respect to the input distribution. Exploring the limitations of current universality results and conditions under which they break remains an interesting direction of research.
@@
  • Jacot et al. [2020] and Bordelon et al. [2020] considered kernel methods for regression and corroborated their findings through experiments on MNIST and Higgs datasets, providing evidence of Gaussian universality.
  • Pandey et al. [2024] explore heavy-tailed data generation using diffusion models. They also observe that this is not possible if one passes Gaussian noise to the models as is usually done and suggest using a Student t-distribution instead. They then demonstrate numerically that their scheme works well for generating the HRRR dataset Dowell et al. [2022].
@@
  • Li et al. [2024b] explore a connection between diffusion models and GMMs from a different point of view. They observe that if the denoisers are over-parametrized, the diffusion models arrive at the GMMs with the means and covariances matching those of the training dataset, but learn to diverge at later stages of training. Our observations imply that even though the distribution of the diffusion-generated images stops being the same as the corresponding GMM after sufficiently many training steps, they still have many properties in common.
@@
  • Concurrent to the submission of this work, Tam and Dunson [2025] published a preprint that establishes a similar Concentration of Measure Property. While their results are valid for any generative model consisting of Lischitz operations, they mainly explore concentration properties for GANs. Our paper conducts extensive numerical experiments with diffusion and dives into the question of bounding the Lipschitz constant of the diffusion process after $N$ steps, which is crucial to ensure that the constants in the concentration inequality can be taken to be independent of $N$. Finally, the second part of Tam and Dunson [2025] considers more abstract settings, such as generative models taking general subgaussian noise as input, while in the second part of our work we proceed to study Gaussian universality for diffusion-generated data.
3 Preliminaries and the theory
We start by defining a key notion needed for our results and then move on to provide an overview of the sampling process of diffusion models. We also prove a universality result for linear models in multiclass classification tasks. We conclude this section by stating our main findings.
3.1 Concentration of Measure Property
We use the following definition of concentration. Informally, it means that the tails of the distribution decay exponentially fast. Note that it corresponds to the Lipschitz Concentration Property for $q=2$ from Seddik et al. [2020].
Definition 1 (CoM).
Given a probability distribution $x\sim\mathbb{P}$ where $\mathbf{x}\in\mathbb{R}^{d}$, we say that $\mathbf{x}$ satisfies the Concentration of Measure Property (CoM) if there exist $C,\sigma>0$ such that for any $L$-Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$ it holds that
$\mathbb{P}\left(|f(\mathbf{x})-\mathbb{E}f(\mathbf{x})|>t\right)\leq Ce^{-(\frac{t}{L\sigma})^{2}}$ (1)
The distributions satisfying CoM arise naturally in many applications and are quite ubiquitous. We appeal to the following proposition proven in Rudelson and Vershynin [2013] :
Proposition 1.
The distribution $\mathbf{x}\sim\mathcal{N}(0,\bm{\Sigma})$ satisfies the CoM property 1. Moreover, the corresponding $C=2$ and $\sigma=\|\bm{\Sigma}^{\frac{1}{2}}\|_{op}$.
If $\bm{\Sigma}=\frac{\mathbf{I}_{d}}{d}$ and $f(\mathbf{x})=\|\mathbf{x}\|_{2}$, then Proposition 1 implies the classical fact that the norm of the normalized standard vector converges to $1$ in probability as $d\to\infty$ because in this case the upper bound of Definition 1 becomes $2e^{-(\frac{t}{\sigma})^{2}}=2e^{-td}\to 0$. However, if $\bm{\Sigma}$ is also normalized as $\text{Tr}(\bm{\Sigma})=1$ but $\|\bm{\Sigma}\|_{op}=\Theta(1)$ (which happens, for instance, if the ordered eigenvalues of $\bm{\Sigma}$ follow the power law $\lambda_{i}=Ci^{-\alpha}$ for some $C>0$ and $\alpha>1$), then the variance of $\mathbf{x}$ does not have to go to zero anymore, but Definition 1 still implies that $\mathbf{x}$ has exponentially decaying tails (to be more precise, $\mathbf{x}$ is a sub-Gaussian random vector–see Definition 3.4.1 in Vershynin [2018]) and in particular cannot be heavy-tailed. The latter was empirically demonstrated via an extensive set of experiments in Pandey et al. [2024] and therefore our results can be considered to be a theoretical validation of the corresponding body of simulations presented in Pandey et al. [2024]. Gaussians are far from the only distributions satisfying CoM; other examples include the strongly log-concave distributions, and the Haar measure-we refer to Section 5.2 in Vershynin [2018] for more examples. The concentration of measure phenomenon has played a key role in the development of many areas such as random functional analysis, compressed sensing, and information theory.
3.2 Diffusion
We provide an overview of diffusion models pertinent to our results in this paper. Given samples $x_{0}\sim q_{0}$ from a high-dimensional distribution in $\mathbb{R}^{d}$, we learn a distribution $p_{\theta}\approx q_{0}$ that allows easy sampling. A trained diffusion model essentially applies a sequence of nonlinear mappings (specifically, denoisers, denoted by $D_{\theta}$) to a white Gaussian input to obtain clean images. Following the formulation in Karras et al. [2022], assuming the distribution of the training to be "delta dirac", the score function can be expressed in terms of the ideal denoiser that minimizes $L_{2}$ error for every noise scale, i.e. $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)=(D_{\theta}(\mathbf{x};\sigma)-\mathbf{x})/\sigma^{2}$. This serves as a heuristic for using $(D_{\theta}(\mathbf{x};\sigma)-\mathbf{x})/\sigma^{2}$ as a surrogate for the score function to run the backward process. In most applications, $D_{\theta}$ is a neural network trained to be a denoiser, typically using a U-Net backbone. The specific denoiser we consider for our experiments is from ADM Dhariwal and Nichol [2021] which uses a modified U-Net backbone with self-attention layers. During training, the network sees multiple noise levels, and learns to denoise the images at many scales. Our analysis and statements in Section 3.5 hold for most of the diffusion models used in practice, as they employ a Lipschitz neural network. In the view of the discussion above, and setting $\sigma(t)=t$, we can represent the sampling process as an iterative procedure of $N$ steps, expressed as:
$\begin{split}&\mathbf{x}_{0}\approx\mathbf{x}^{(N)}=\mathcal{R}^{(N-1)}_{D_{\theta}}\Biggl{(}\mathcal{R}^{(N-2)}_{D_{\theta}}\biggl{(}\\ &\Bigl{(}\ldots\mathcal{R}^{(0)}_{D_{\theta}}\bigl{(}\mathbf{x}^{(0)},t_{0:1}\bigr{)}\ldots\Bigr{)},t_{N-2:N-1}\biggr{)},t_{N-1:N}\Biggr{)}\end{split}$ (2)
Figure 1: High-level overview of the sampling process in the diffusion models
  $\bullet$ Prob Flow ODE
1 Sample $\mathbf{x}^{(0)}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I})$
2 for  $i\in\{0,\ldots,N-1\}$  do
3       Sample $\bm{\epsilon}\sim\mathcal{N}(0,S_{\text{noise}}^{2}\mathbf{I})$
        $\bullet$ Inject Noise
       $h\leftarrow t_{i+1}-t_{i}(1+\gamma)$ ;
        $\bullet$ Step size
       $\mathbf{x}^{(i+1)}\leftarrow\hat{\mathbf{x}}^{(i)}+hf(\hat{\mathbf{x}}^{(i)},t_{i}(1+\gamma))$ ;
        $\bullet$ Euler step
4       if $t_{i+1}\neq 0$ then
             $\mathbf{x}^{(i+1)}\leftarrow\hat{\mathbf{x}}^{(i)}+\frac{h}{2}\left(f(\hat{\mathbf{x}}^{(i)},t_{i}(1+\gamma))+f(\mathbf{x}^{(i+1)},t_{i+1})\right)$ ;
              $\bullet$ Second-order correction
5
6
return $\mathbf{x}^{(N)}$
Algorithm 1 EDM Stochastic Sampler Karras et al. [2022]
In this work, we focus on the framework considered in Karras et al. [2022] and we observe that in summary, $\mathbf{x}^{(i+1)}\leftarrow\mathcal{R}_{D_{\theta}}^{(i)}(\mathbf{x}^{(i)},t_{i:i+1})$, with
$\mathcal{R}_{D_{\theta}}^{(i)}(\mathbf{x}^{(i)},t_{i:i+1}):=\hat{\mathbf{x}}^{(i)}$
$+\frac{h}{2t_{i+1}}\left[\hat{\mathbf{x}}^{(i)}+(h+t_{i+1})d_{i}-\underbrace{D_{\theta}(\hat{\mathbf{x}}^{(i)}+hd_{i},t_{i+1})}_{\text{Denoiserafter Euler step}}\right]$
Where $d_{i}:=f(\hat{\mathbf{x}}^{(i)},t_{i}(1+\gamma))$ is as defined in line 1 of Algorithm 1 and $\gamma$ is a hyperparameter controlling the amount of additional injected noise whose scale is determined by the $S_{\text{noise}}$ hyperparameter. And $\hat{\mathbf{x}}^{(i)}$ is the current image with the added noise. Formally, we would like to claim that the distribution of the output $\mathbf{x}^{(N)}$ satisfies CoM, and we visualize the evolution of the norms of these quantities through the sampling process in Figure 4 to further illuminate our argument about the $1$-Lipschitznes of the generative process.
@@ 3.3 Classification and Gaussian Universality
We cover Gaussian universality in the context of linear multiclass classification following the framework described in Ghane et al. [2024] and extend it to an arbitrary number of classes. As we will see, most known Gaussian universality results operate in an idealized setting that does not appear to be applicable to the covariance matrices estimated from the diffusion-generated images (Figure 9). Nevertheless, we observe empirically that universality holds in the latter setting as well, hence raising a challenge of relaxing the assumptions of the existing universality results to make them more practical. We outline the corresponding notation and challenge below.
@@
    $\mathbf{X}=\begin{pmatrix}\mathbf{x}_{1}^{T}\\ \mathbf{x}_{2}^{T}\\ \vdots\\ \mathbf{x}_{n}^{T}\end{pmatrix},\quad\mathbf{Y}=\begin{pmatrix}\mathbf{y}_{1}^{T}\\ \mathbf{y}_{2}^{T}\\ \vdots\\ \mathbf{y}_{n}^{T}\end{pmatrix}$
    Without loss of generality, we can rearrange the rows of $\mathbf{X}$ to group samples from the same class together.
@@
    To train for $\mathbf{W}$, we minimize $\|\mathbf{Y}-\mathbf{X}\mathbf{W}\|_{F}^{2}$ by running SGD with a constant stepsize. By the implicit bias property of SGD Gunasekar et al. [2018], Azizan and Hassibi [2018] for linear models, we observe that the iterations of SGD initialized from some $\mathbf{W}_{0}$ converge to the optimal solution of the following optimization problem
    $\min_{\mathbf{W}\in\mathbb{R}^{d\times k}}\|\mathbf{W}-\mathbf{W}_{0}\|_{F}^{2}$ (3)
    $s.t\quad\mathbf{X}\mathbf{W}=\mathbf{Y}$ (4)
@@
    $\min_{\tilde{\mathbf{W}}\in\mathbb{R}^{d\times k}}\|\tilde{\mathbf{W}}-\mathbf{W}_{0}\|_{F}^{2}$
    $s.t\quad\mathbf{G}\tilde{\mathbf{W}}=\mathbf{Y}$
@@
The following holds asymptotically under Assumptions 1 :
$\Biggl{|}\sum_{i=1}^{k}\theta_{i}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell}^{T}\mathbf{x}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}$
$-\sum_{i=1}^{k}\theta_{i}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\tilde{\mathbf{W}}_{\ell}^{T}\mathbf{g}\Bigl{|}\mathbf{g}\sim\mathcal{N}(\bm{\mu}_{i},\bm{\Sigma}_{i}))\Biggr{|}\to 0$
@@
  • $s_{\min}(\mathbf{X}\mathbf{X}^{T})=\Omega(1)$ with high probability where $s_{\min}(.)$ is the smallest singular value.

### v3

# Gaussian Universality for Diffusion Models

@@ Abstract
We investigate Gaussian Universality for data distributions generated via diffusion models. By Gaussian Universality we mean that the test error of a generalized linear model $f(\mathbf{W})$ trained for a classification task on the diffusion data matches the test error of $f(\mathbf{W})$ trained on the Gaussian Mixture with matching means and covariances per class. In other words, the test error depends only on the first and second order statistics of the diffusion-generated data in the linear setting. As a corollary, the analysis of the test error for linear classifiers can be reduced to Gaussian data from diffusion-generated data. Analysing the performance of models trained on synthetic data is a pertinent problem due to the surge of methods such as Sehwag et al. (2024). Moreover, we show that, for any $1$- Lipschitz scalar function $\phi$, $\phi(\mathbf{x})$ is close to $\mathbb{E}\phi(\mathbf{x})$ with high probability for $\mathbf{x}$ sampled from the conditional diffusion model corresponding to each class. Finally, we note that current approaches for proving universality do not apply to diffusion-generated data as the covariance matrices of the data tend to have vanishing minimum singular values, contrary to the assumption made in the literature. This leaves extending previous mathematical universality results as an intriguing open question.
@@ 1 Introduction
A remarkable contribution of deep learning is the advent of generative models for image and video generation. Diffusion-based generative models Sohl-Dickstein et al. (2015), Song and Ermon (2019),Ho et al. (2020),Song et al. (2020),Dhariwal and Nichol (2021),Song et al. (2021),Kingma et al. (2021), Karras et al. (2022), in particular, have enjoyed tremendous success in vision [LDM Rombach et al. (2022), audio [Diffwave Kong et al. (2020)] and text generation [D3PM Austin et al. (2021)]. For an overview of diffusion models and their applications, we refer to the surveys Croitoru et al. (2023) and Yang et al. (2023).
Despite significant progress in training methods, network architecture design, and hyperparameter tuning, there has been relatively little work done on understanding rigorous mathematical properties of the data generated by diffusion models. Through theory and experiments, we argue that images generated by conventional diffusion models satisfy a form of Gaussian Universality.
Gaussian Universality is an overloaded term. It often means that certain characteristics of data coming from $k$ classes distributed as $\mathbb{P}_{1},\dots,\mathbb{P}_{k}$ with means $\bm{\mu}_{1},\dots,\bm{\mu}_{k}$ and covariance matrices $\bm{\Sigma}_{1},\dots,\bm{\Sigma}_{k}$ remain unchanged when replaced by data coming from the corresponding Gaussians $\mathcal{N}(\bm{\mu}_{1},\bm{\Sigma}_{1}),\dots,\mathcal{N}(\bm{\mu}_{k},\bm{\Sigma}_{k})$. Informally, this means that a data representation satisfying Gaussian Universality behaves similarly to a Gaussian Mixture Model with matching means and covariances per class.
Gaussian Universality has been observed in the empirical distribution of eigenvalues of Gram matrices constructed from real-world data Seddik et al. (2020); Levi and Oz (2023). While this phenomenon is compelling, we focus on Gaussian Universality of the test error, which we consider a key characteristic in machine learning. Specifically, we compare the generalization error of generalized linear models trained on diffusion-generated data to those trained on Gaussian Mixtures with matching means and covariances, observing a close match.
As a technical step towards elucidating this phenomenon, we argue that, when the reverse process is a contraction, one can establish an Approximate Concentration of Measure phenomenon for the distribution of the output, which is a common assumption in theoretical works on Gaussian Universality of the test error.
@@
  • We show that distributions that can be generated via diffusion models satisfy an Approximate Concentration of Measure Property 1. Similar properties are required in most apporaches to proving universality.
  • As discussed in Goldblum et al. (2023) and Nakkiran (2021), one of the reasons we do not have a practical theory of deep learning is the lack of clean mathematical models that describe real-world data. In view of the universality results outlined in the presented work, combined with the empirical observations showing that diffusion-generated distributions can approximate real-world data well, we suggest that it suffices to consider GMMs. The analysis of performance of models trained on GMMs has been a topic of active research recently; see, e.g. Thrampoulidis et al. (2020); Loureiro et al. (2021a).
Finally, we note that while this paper primarily assesses Gaussian Universality of test error for diffusion models trained on real-world data, the theoretical section is self-contained and presents all necessary mathematical formulations, results, and references.
@@ 2 Related Works
  • In an emerging line of work, many papers have analyzed the output distributions and convergence of diffusion models through the lens of Langevin dynamics. Chen et al. (2022) show that denoising diffusion probabilistic models (DDPM) and critically damped Langevin Dynamics (CLD) can efficiently sample from any arbitrary distribution, assuming accurate score estimates - an assumption central to many works in this area. While among the first works to provide quantitative polynomial bounds on convergence, the high-dimensional nature of the problem means that estimating the score function may be practically impossible. Furthermore, it is infeasible to verify the validity of this assumption, as we do not have access to the true score function. And, as evident from the bounds of Mousavi-Hosseini et al. (2023), generating heavy-tailed distributions using Langevin dynamics initialized from the Gaussian distribution is intractable in practice as one needs to run the Langevin dynamics for an exponential number of steps. We refer to Li et al. (2024a) for a brief overview of the existing works on the convergence theory of diffusion models.
@@
  • Seddik et al. (2020) show a form of equivalence between representations generated from Generative Adversarial Networks (GANs) and from GMMs. They considered the Gram matrix of pre-trained classifier representations of the GAN-generated images and show that asymptotically, it possesses the same distribution of eigenvalues as the Gram matrix of Gaussian samples with matching first and second moments.
@@
  • Loureiro et al. (2021a) investigated the generalization error of linear models for binary classification with logistic loss and $\ell_{2}$ regularization. On MNIST and Fashion MNIST, they observed a close match between the real images and the corresponding GMM for the linear models and in the feature map of a two-layer network. Loureiro et al. (2021b) considered a student-teacher model and verified universality for the aforementioned datasets via kernel ridge regression. They also explored the output of a deep convolutional GAN (dcGAN), labeling it using a three-layer teacher network. Using logistic regression for classification illustrated a close match with GMMs on GAN-generated data, but a deviation from real CIFAR10 images. Goldt et al. (2022) analyzed the generalization error of Random Features logistic regression using the Gaussian Equivalence property and corroborated their results using images generated by a dcGAN trained on CIFAR100. Pesce et al. (2023) studied the student-teacher model for classification and empirically demonstrated the universality of the double descent phenomenon for MNIST and Fashion MNIST. They preprocessed these datasets using a random feature map, with labels generated by a random teacher, for ridge regression and logistic classification. However, they also observed that the universality of the test error fails to hold while using CIFAR10 without preprocessing with either random feature maps, wavelet scattering, or Hadamard orthogonal projection. Dandi et al. (2024) observed that the data distributions generated by conditional GANs trained on Fashion MNIST exhibit Gaussian universality of the test error for generalized linear models. Gerace et al. (2024) considered mixture distributions with random labels and demonstrated universality of test error of the generalized linear models. The universality part of our work can be considered as an exploration of the same phenomenon for conditional diffusion models trained on significantly larger image datasets.
@@
  • Refinetti et al. (2023) show that SGD learns higher moments of the data as the training continues which exhibited nonuniversality of the test error with respect to the input distribution. Exploring the limitations of current universality results and conditions under which they break remains an interesting direction of research.
@@
  • Jacot et al. (2020) and Bordelon et al. (2020) considered kernel methods for regression and corroborated their findings through experiments on MNIST and Higgs datasets, providing evidence of Gaussian universality.
@@
  • Li et al. (2024b) explore a connection between diffusion models and GMMs from a different point of view. They observe that if the denoisers are over-parametrized, the diffusion models arrive at the GMMs with the means and covariances matching those of the training dataset, but learn to diverge at later stages of training. Our observations imply that even though the distribution of the diffusion-generated images stops being the same as the corresponding GMM after sufficiently many training steps, they still have properties in common.
  • Levi and Oz (2023) investigated the spectrum of the empirical Gram and covariance matrices of various real world image datasets by centering the images. They demonstrated that the eigendistribution and the level-spacing distribution of the empirical covariance matrix of real data could be closely captured with that of a Wishart matrix whose covariance displays a Toeplitz structure with eigenvalues following a power-law spectrum. Albeit not the main focus of our work, we show that the spectrum of the covariance matrix for each class of data generated by a conventional conditional diffusion model also displays a power-law behavior. Our experiment setting is different in two key aspects. First, by not centering the images, we retain the mean of each class in the data. Second, the empirical covariance is compute per class, rather than over the entire dataset.
@@
  • Concurrent to the submission of this work, Tam and Dunson (2025) established a similar Concentration of Measure Property. While their results are valid for any generative model consisting of Lipschitz operations, they mainly explore concentration properties for GANs. Our paper conducts extensive numerical experiments with diffusion and dives into the question of bounding the Lipschitz constant of the diffusion process after $N$ steps, which is crucial to ensure that the constants in the concentration inequality can be taken to be independent of $N$. Finally, the second part of Tam and Dunson (2025) considers more abstract settings, such as generative models taking general subgaussian noise as input, while in the second part of our work we study Gaussian universality for diffusion-generated data.
3 Main Results
We start by defining the linear multiclass classification problem and we prove a universality result for such models and state our first empirical observation. We then proceed to provide insights and explanations for why this empirical may hold based on a notion of approximate concentration of measure which requires an analysis of the sampling process of diffusion models (Appendix A).
@@ 3.1 Classification and Gaussian Universality
We cover Gaussian universality in the context of linear multiclass classification following the framework described in Ghane et al. (2024) and extend it to an arbitrary number of classes. As we will see, most known Gaussian universality results operate in an idealized setting that does not appear to be applicable to the covariance matrices estimated from the diffusion-generated images (Figure 4). Nevertheless, we observe empirically that universality holds in the latter setting as well, hence raising a challenge of relaxing the assumptions of the existing universality results to make them more practical. We outline the corresponding notation and challenge below.
@@
    $\mathbf{X}^{T}:=\begin{pmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\ldots&\mathbf{x}_{n}\end{pmatrix},\quad\mathbf{Y}^{T}:=\begin{pmatrix}\mathbf{y}_{1}&\mathbf{y}_{2}&\ldots&\mathbf{y}_{n}\end{pmatrix}$
    Without loss of generality, we can rearrange rows of $\mathbf{X}$ to group together samples by class.
@@
  • To train for $\mathbf{W}$, we minimize $\|\mathbf{Y}-\mathbf{X}\mathbf{W}\|_{F}^{2}$ by running SGD with a constant stepsize. By the implicit bias property of SGD Gunasekar et al. (2018), Azizan and Hassibi (2018) for linear models, we observe that the iterations of SGD initialized from some $\mathbf{W}_{0}$ converge to the optimal solution of the following optimization problem
    $\min_{\mathbf{W}\in\mathbb{R}^{d\times k}}\|\mathbf{W}-\mathbf{W}_{0}\|_{F}^{2}\quad s.t\quad\mathbf{X}\mathbf{W}=\mathbf{Y}$ (1)
@@
    $\min_{\tilde{\mathbf{W}}\in\mathbb{R}^{d\times k}}\|\tilde{\mathbf{W}}-\mathbf{W}_{0}\|_{F}^{2}\quad s.t\quad\mathbf{G}\tilde{\mathbf{W}}=\mathbf{Y}$
@@
The following holds asymptotically in the proportional regime $\frac{d}{n}\rightarrow\delta>1$ under Assumptions 1 :
$\Biggl|\sum_{i=1}^{k}\theta_{i}\mathbb{P}\Bigl(i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{w}_{\ell}^{T}\mathbf{x}\Bigl|\mathbf{x}\sim\mathbb{P}_{i}\Bigr)-\sum_{i=1}^{k}\theta_{i}\mathbb{P}\Bigl(i\neq\operatorname*{arg\,max}_{\ell\in[k]}\tilde{\mathbf{w}}_{\ell}^{T}\mathbf{g}\Bigl|\mathbf{g}\sim\mathcal{N}(\bm{\mu}_{i},\bm{\Sigma}_{i})\Bigr)\Biggr|\to 0$
@@
  • $s_{\min}(\mathbf{X}\mathbf{X}^{T})=\Omega(1)$ with high probability where $s_{\min}(\cdot)$ is the smallest singular value.